solve using substitution method
4=3x-5
4x+y=23

Answer by Alan3354(69443)

solve using substitution method
4=3x-5
4x+y=23
----------
4=3x-5
3x=9
x = 3
--------
4x+y = 23
12+y = 23
y = 11